Show that the triangle with vertices A(2,2), B(-1,1) and C (1,-1) is equilateral under the taxicab metric, but not under the euclidean metric
1
Expert's answer
2020-10-13T19:34:05-0400


In the taxicab metric the length of edges:

"AB=|A_{x}-B_{x}|+|A_{y}-B_{y}|="

"=|2-(-1)|+|2-1|=4" or "AB=2\Delta y+3 (2\Delta x)=2\cdot0.5+6\cdot0.5=4;"

"BC=|B_{x}-C_{x}|+|B_{y}-C_{y}|="

"=|-1-1|+|1-(-1)|=4;"

"AC=|A_{x}-C_{x}|+|A_{y}-C_{y}|="

"=|2-1|+|2-(-1)|=4" .

"AB=BC=AC=4" therefore the triangle is equilateral under the taxicab metric.


In the euclidean metric the length of edges:

"AB=\sqrt{(A_{x}-B_{x})^2+(A_{y}-B_{y})^2}="

"=\sqrt{(2-(-1))^2+(2-1)^2}=\sqrt{3^2+1^2}="

"=\sqrt{10}\approx3.16;"

"BC=\sqrt{(B_{x}-C_{x})^2+(B_{y}-C_{y})^2}="

"=\sqrt{(-1-1)^2+(1-(-1))^2}=\sqrt{(-2)^2+2^2}="

"=\sqrt{8}\approx2.83;"

"AC=\sqrt{(A_{x}-C_{x})^2+(A_{y}-C_{y})^2}="

"=\sqrt{(2-1)^2+(2-(-1))^2}=\sqrt{1^2+3^2}="

"=\sqrt{10}\approx3.16" .

"AB=AC\neq BC" therefore the triangle is not equilateral under the euclidean metric.